A simple harmonic oscillator of frequency 1Hz has a phase of 1 radian.By how much should the origin be shifted in time so as to make the phase of the oscillator vanish. (time in seconds (s)).
Ans: -1/2π
1
Expert's answer
2020-09-22T15:36:23-0400

A phase of a simple harmonic oscillator is given by the following expression:


"\varphi = \omega t"

where "\omega = 2\pi f = 2\pi\cdot 1Hz = 2\pi" is the angular frequency, "t" is the time. If "\varphi = 1", then


"t = \dfrac{\varphi}{\omega} = \dfrac{1}{\omega} = \dfrac{1}{2\pi}"

Thus, if we shift time by "-1/2\pi", we get rid of the phase ("\varphi = 0" ).


Answer. "t = \dfrac{1}{2\pi}" s.